OpenStudy (anonymous):

find volume if r = 4 and the slant and height of the cone is 5 helpp

OpenStudy (watchmath):

Using Pythagorean we have \(h^2+4^2=5^2\) \(h^2=9\) \(h=3\) Then the volume is \(V=\frac{1}{3}\pi r^2h=\frac{1}{3}\pi (4)^2(3)=16\pi\)